# Problem-informed Graphical Quantum Generative Learning

### v1

@@ Abstract
Leveraging the intrinsic probabilistic nature of quantum systems, generative quantum machine learning (QML) offers the potential to outperform classical learning models. Current generative QML algorithms mostly rely on general-purpose models that, while being very expressive, face several training challenges. A potential way to address these setbacks involves constructing problem-informed models capable of more efficient training on structured problems. In particular, probabilistic graphical models provide a flexible framework for representing structure in generative learning problems and can thus be exploited to incorporate inductive bias in QML algorithms. In this work, we propose a problem-informed quantum circuit Born machine Ansatz for learning the joint probability distribution of random variables, with independence relations efficiently represented by a Markov network (MN). We further demonstrate the applicability of the MN framework in constructing generative learning benchmarks and compare our model’s performance to previous designs, showing it outperforms problem-agnostic circuits. Based on a preliminary analysis of trainability, we narrow down the class of MNs to those exhibiting favorable trainability properties. Finally, we discuss the potential of our model to offer quantum advantage in the context of generative learning.
@@ I Introduction
While their high expressivity makes general-purpose QML models very powerful, they also pose several challenges. Contrary to classical neural networks, variational quantum circuits are much more affected by trainability issues, such as barren plateaus and poor local minima [18, 19, 20, 21, 22]. Furthermore, the no-free-lunch theorem [23] also translates to QML, suggesting, that these problem-agnostic models, like hardware-efficient Ansätze, have poor average performance [24, 25]. The reason behind these barriers can be seen as the lack of sufficient inductive bias, i.e., assumptions about the data that could be encoded into the learning framework. Consequently, a potential way of dealing with them is by constructing problem-informed models, that can be trained more efficiently for structured problems [26, 27, 28, 29].
@@
While there are several excellent works concerning the quantum circuit implementation of BNs [33, 34, 35], MNs are not well-studied in the context of QML. In this work, we investigate the applicability of the framework provided by MNs to generative QML with classical data. We propose a problem-informed model, that aims to learn the distribution over random variables, where the independence relations are efficiently represented by a MN. As opposed to previous problem-agnostic QCBMs, this Ansatz can capture higher order correlations between the corresponding random variables and potentially reduce the number of trainable parameters, while also increasing performance. While this construction relies on the knowledge of the MN structure, that can be hard to infer from data, the graph representation is readily available in various application domains. We argue that this new model class has the potential to demonstrate quantum advantage, since it contains the class of QAOA circuits [36], that were shown to produce classically hard probability distributions [7, 37]. Besides model design, we also show the potential of the PGM framework to construct benchmarks for generative QML models, where the problem complexity can be fine-tuned in multiple factors. We perform numerical experiments based on this benchmark proposal against both problem-agnostic and BN-based QML models. Finally, we present a preliminary analysis of trainability and define a class of efficient graphical representations, that have higher potential in this context.
@@ II.1 Probabilistic Graphical Modelling
Here we give a brief description of the mathematical framework of the two main families of PGMs and their connections. Since any higher order PGM can be embedded into PGMs over binary random variables, we focus our attention on the latter. For an extensive study of PGMs, we refer the reader to [30].
@@ II.1.2 Markov networks
Markov networks, Markov random fields or undirected graphical models are defined over general undirected graphs to represent a set of random variables having the Markov property. In general, the global Markov property applies, which states, that any two subsets of variables are conditionally independent given a separating subset. As opposed to BNs, where the factors are straightforward to comprehend, here they cannot be interpreted directly. However, we can view a factor as describing “compatibilities” between different values of the variables in the corresponding subset. A factor here describes a general purpose function $\phi:\mathrm{Val}(\mathbf{D})\rightarrow\mathbb{R}$, where $\mathrm{Val}(\mathbf{D})$ denotes all possible joint states of a set of random variables $\mathbf{D}$. Each factor corresponds to a clique in the graph, however, the usual graphical representation does not make it clear, whether the joint probability distribution factorizes according to the maximal cliques or the subsets thereof.
@@
$\begin{split}P_{\Phi}(X_{1},\dots,X_{N})&=\frac{1}{Z}\tilde{P}_{\Phi}(X_{1},\dots,X_{N})\\ &=\frac{1}{Z}\phi_{1}(\mathbf{D}_{1})\times\phi_{2}(\mathbf{D}_{2})\times\dots\times\phi_{m}(\mathbf{D}_{m}),\end{split}$
@@
For the MN in Fig. 2b, with independence relations $(A\perp D\;|\;C)$, $(B\perp D\;|\;C)$ and assuming maximal clique factorisation, the joint probability distribution takes the form $P(A,B,C,D)=\frac{1}{Z}\phi_{1}(A,B,C)\phi_{2}(C,D)$. Here $\phi_{1}(A,B,C)$ and $\phi_{2}(C,D)$ denote the two maximal cliques of the graph.
@@
where $\epsilon_{i}(\mathbf{D}_{i})=-\ln{\phi_{i}(\mathbf{D}_{i})}$ is called the energy function. This parametrization of MNs is usually referred to as the log-linear model. We use this representation as inspiration for constructing our MN-based QML model.
@@ II.1.3 Connection between BNs and MNs
The transformation from BNs to MNs, called moralization, follows a simple rule: if two nodes are connected by a directed edge in the BN graph or they are both parents of at least one node, then they are connected by an undirected edge in the MN graph. Having the undiercted graph, we can assign a factor to each resulting clique, to obtain a MN. This procedure usually leads to a PGM with more parameters, that can lead to a longer training. Directed graphs, in which parent nodes do not share common children are called moral. Consequently, BNs that are defined over moral graphs, can be converted to MNs without introducing additional edges and parameters.
Turning MNs into BNs is a more difficult task, both conceptually and computationally. First, the undirencted graph needs to be triangulated, meaning that we introduce chords in cycles of $4$ or more and repeat this process until there are no such cycles left. Chords connect nodes, that are in the cycle, but are not already connected. This transformation usually leads to the introduction of a much larger number of edges than in the previous case. Finally, the undirected edges have to be turned into directed ones in an acyclic manner. Since chordal graphs are also moral, PGMs that are defined over chordal graphs represent a class of graphical models, that can be treated as either BNs or MNs and the conversion between the two is straightforward.
@@ II.1.4 Generative learning in PGMs
In MNs, the use of a normalizing constant couples the parameters across the whole network, which prevents us from decomposing the problem and estimating local groups of parameters separately. One of the computational ramifications of this global coupling is, that not even the parameter estimation with complete data can be solved in closed form (except for some special cases, that can essentially be reformulated as BNs [40]). This makes the use of iterative methods, such as gradient descent, unavoidable. Luckily, the likelihood objective is convex, meaning that these methods are guaranteed to converge. However, they come with the disadvantage of having to run inference in each step to calculate the gradients and inference in MNs is #P-complete in general, which makes our distribution learning task formulated above fairly expensive or even intractable classically [41].
@@ II.2.1 General framework of QCBMs
QCBMs, introduced in [11], are paradigmatic generative QML models, that naturally inherit the Born rule and thus can be used to generate tunable and discrete probability distributions that approximate a target distribution. In this section, we review the general components of variational quantum algorithms in the context of QCBMs.
One of the key building blocks is the Ansatz. The Ansatz refers to a parametrized family of quantum circuits used as a hypothesis or approximation for solving the problem in question by iteratively optimizing the parameters to match the desired target behavior or to minimize a cost function. We can differentiate problem-informed and problem-agnostic Ansätze.
Figure 3: General structure of the quantum circuit Ising Born machine adapted from [13]. The circuit starts with Hadamard gates on each qubit, then a parametrized $n$-qubit unitary is applied with tunable parameters $\boldsymbol{\alpha}$, finally, a general parametrized $U_{f}$ gate acts on each qubit separately. The measurement is taken in the computational basis.
@@
$U_{f}(\mathbf{\Gamma},\mathbf{\Delta},\mathbf{\Sigma})=\exp{\left(i\sum\limits_{k=1}^{n}(\Gamma_{k}X_{k}+\Delta_{k}Y_{k}+\Sigma_{k}Z_{k})\right)}.$ (4)
@@
There are multiple ways to characterize the distance of two distributions. The TV distance, shown in (2), is a good benchmark, but it is not feasible as a cost function. One of the most common metrics used as a cost function in generative modelling is the Kullback-Leibler (KL) divergence
@@
Unfortunately, not even this function is optimal for data-driven training, as it requires a large number of training samples and it was shown not to be trainable for larger-scale QCBMs [42]. In [12] a more efficient cost function was introduced for gradient-based training of QCBMs. The idea was to compare the distance of the samples drawn from the target and the model distribution in a kernel feature space. This loss function is called the squared maximum mean discrepancy (MMD):
@@
Having a family of parametrized circuits and a cost function to characterize the distance between the circuit output and the target distribution, the final building block is the optimizer, that adjusts the circuit parameters given the cost function. Both gradient-based and gradient-free optimizers have been used to train QCBMs [12, 43]. We concentrate on gradient-based optimization methods, but since our approach concentrates on Ansatz-design, it is compatible with other optimizers as well.
@@ II.2.2 Existing quantum circuit adaptations of PGMs
Liu and Wang [12] made the first explicit connection between the paradigmatic QCBM model and PGMs. They propose a framework, in which they first construct the Chow-Liu tree of a dataset [44] based on the mutual information between all pairs of bits for training samples. Having this tree graph, they propose a QCBM Ansatz, in which, the connectivity pattern of the $CNOT$ gates respects the graph structure. The Chow-Liu tree offers an effective approach for creating a second-order product approximation of a joint probability distribution. The corresponding graph represents a BN, that can also be regarded as a pairwise MN, but being a second-order approximation, it fails to detect higher order correlations. Another explicit connection between BNs and QCBMs was formulated in [45], where the authors proposed a framework, that utilizes QCBMs for variational inference in PGMs.
@@
Besides the framework of mostly general-purpose QCBMs, there have been several attempts to implement PGMs on a quantum computer. Bayesian networks have an equivalent formulation in the computational basis measurements of a class of quantum circuits known as Bayesian quantum circuits (BQCs) [33]. These are defined such that the probability distribution they sample from, by measuring the given qubits in the computational basis, corresponds to the distribution defined by the corresponding BN. BQCs are implemented with uniformly controlled gates, that can be decomposed into one-qubit rotations and $CNOT$ gates [34, 46]. By definition, these circuits obey certain rules in accordance with the directed, acyclic nature of the underlying graph.
@@
In [35] the authors introduced a minimal extension to BQCs and presented unconditional proof of separation in the expressive power of BNs and the corresponding basis-enhanced BQCs (BBQCs). They showed that by letting the final measurement be in any local basis, this separation appears, that can be associated with quantum nonlocality and contextuality. They also pointed out, that both BQCs and their basis-enhanced versions can be efficiently simulated with classical tensor network methods, when the graphs are sparse enough.
@@
The literature on the quantum circuit implementation of Markov networks is more scarce. In Ref. [47] the authors identified a novel embedding technique of MNs into unitary operators that relies on their log-linear representation. They construct a Hamiltonian composed of Pauli-$Z$ terms and give a quantum algorithm, that implements the exponential of this Hamiltonian, meaning, that measuring the output qubits of the corresponding quantum circuit is equivalent to sampling the corresponding MN. The circuit, that implements this exponentiation, uses a special point-wise polynomial approximation and real part extraction, that might fail. For this reason, they have to measure ancillary qubits in order to determine whether this extraction was successful and start everything over if not. Consequently, the success probability decreases exponentially with the number of maximal cliques. This can of course be amplified with quantum singular value transformation [48], but that further increases the required resources. Since Boltzmann machines form a subclass of pairwise MNs, their quantum circuit implementations can also be regarded as an adaptations of PGMs [17]. However, these models are quite restricted compared to general MNs, since they only consider pairwise correlations, usually in a bipartite manner.
@@ III Quantum Circuit Markov Random Fields
In this section, we present our results, starting with the definition of our QML model, proposed for generative learning in MNs. We then introduce our novel benchmark proposal and compare our model to both problem-agnostic QCIBMs and BBQCs through a series of numerical experiments. As a a preliminary analysis of trainability, we investigate the scaling of the cost function variance for different types of graphs. Finally, we present our argument for a potential quantum advantage of our MN-based model class.

### v2

@@ Abstract
Leveraging the intrinsic probabilistic nature of quantum systems, generative quantum machine learning (QML) offers the potential to outperform classical learning models. Current generative QML algorithms mostly rely on general-purpose models that, while being very expressive, face several training challenges. One potential way to address these setbacks is by constructing problem-informed models that are capable of more efficient training on structured problems. In particular, probabilistic graphical models provide a flexible framework for representing structure in generative learning problems and can thus be exploited to incorporate inductive bias into QML algorithms. In this work, we propose a problem-informed quantum circuit Born machine Ansatz for learning the joint probability distribution of random variables, with independence relations efficiently represented by a Markov network (MN). We further demonstrate the applicability of the MN framework in constructing generative learning benchmarks and compare our model’s performance to previous designs, showing that it outperforms problem-agnostic circuits. Based on a preliminary analysis of trainability, we narrow down the class of MNs to those exhibiting favourable trainability properties. Finally, we discuss the potential of our model to offer quantum advantage in the context of generative learning.
@@ I Introduction
While their high expressivity makes general-purpose QML models very powerful, they also pose several challenges. Contrary to classical neural networks, variational quantum circuits are much more affected by trainability issues, such as barren plateaus and poor local minima [18, 19, 20, 21, 22]. Furthermore, the no-free-lunch theorem [23] also translates to QML, suggesting, that these problem-agnostic models, like hardware-efficient Ansätze, have poor average performance [24, 25]. The reason behind these barriers can be seen as the lack of sufficient inductive bias, i.e., assumptions about the data that could be encoded into the learning framework. Consequently, a potential way of dealing with them is by constructing problem-informed models, that can be trained more efficiently for structured problems [26, 27, 28, 29, 30].
@@
While there are several excellent works concerning the quantum circuit implementation of BNs [34, 35, 36], MNs are not well-studied in the context of QML. In this work, we investigate the applicability of the framework provided by MNs to generative QML with classical data. We propose a problem-informed model, that aims to learn the distribution over random variables, where the independence relations are efficiently represented by a MN. As opposed to previous problem-agnostic QCBMs, this Ansatz can capture higher-order correlations between the corresponding random variables and potentially reduce the number of trainable parameters, while also increasing performance. Throughout the paper, by higher-order correlations, we refer to correlations, that cannot be described by fewer-body (e.g., pairwise) relationships. While this construction relies on the knowledge of the MN structure, that can be hard to infer from data, the graph representation is readily available in various application domains. We argue that this new model class has the potential to demonstrate quantum advantage, since it contains the class of QAOA circuits [37], that were shown to produce classically hard probability distributions [7, 38]. Besides model design, we also show the potential of the PGM framework to construct benchmarks for generative QML models, where the problem complexity can be fine-tuned in multiple factors. We perform numerical experiments based on this benchmark proposal against both problem-agnostic and BN-based QML models. Finally, we present a preliminary analysis of trainability and define a class of efficient graphical representations, that have higher potential in this context.
@@ II.1 Probabilistic Graphical Modelling
Here we give a brief description of the mathematical framework of the two main families of PGMs and their connections. Since any higher-order PGM can be embedded into PGMs over binary random variables, we focus our attention on the latter. For an extensive study of PGMs, we refer the reader to Ref. [31].
@@ II.1.2 Markov networks
Markov networks, Markov random fields or undirected graphical models are defined over general undirected graphs to represent a set of random variables having the Markov property. In general, the global Markov property applies, which states, that any two subsets of variables are conditionally independent given a separating subset. As opposed to BNs, where the factors are straightforward to comprehend, here they cannot be interpreted directly. However, we can view a factor as describing “compatibilities” between different values of the variables in the corresponding subset. A factor here describes a general purpose function $\phi:\mathrm{Val}(\mathbf{D})\rightarrow\mathbb{R}$, where $\mathrm{Val}(\mathbf{D})$ denotes all possible joint states of a set of random variables $\mathbf{D}$. Each factor corresponds to a clique in the graph, however, the usual graphical representation does not make it clear, whether the joint probability distribution only factorizes according to the maximal cliques or also the subsets thereof. However, the maximal clique factorization is the most general for a given graph, therefore, in general, we work with this assumption.
@@
$\begin{split}P_{\Phi}(X_{1},\dots,X_{N})&=\frac{1}{Z}\tilde{P}_{\Phi}(X_{1},\dots,X_{N})\\ &=\frac{1}{Z}\phi_{1}(\mathbf{D}_{1})\times\phi_{2}(\mathbf{D}_{2})\times\dots\times\phi_{K}(\mathbf{D}_{K}),\end{split}$
@@
For the MN in Fig. 2b, with independence relations $(A\perp D\;|\;C)$, $(B\perp D\;|\;C)$ and assuming maximal clique factorisation, the joint probability distribution, according to Def. 2, takes the form $P(A,B,C,D)=\frac{1}{Z}\phi_{1}(A,B,C)\phi_{2}(C,D)$. Here $\phi_{1}(A,B,C)$ and $\phi_{2}(C,D)$ denote the two maximal cliques of the graph.
@@
where $\epsilon_{i}(\mathbf{D}_{i})=-\ln{\phi_{i}(\mathbf{D}_{i})}$ ) is called the energy function. This parametrization of MNs is usually referred to as the log-linear model. We use this representation as inspiration for constructing our MN-based QML model.
@@ II.1.3 Connection between BNs and MNs
The transformation from BNs to MNs, called moralization, follows a simple rule: if two nodes are connected by a directed edge in the BN graph, or they are both parents of at least one node, then they are connected by an undirected edge in the MN graph. Having the undirected graph, we can assign a factor to each resulting clique, to obtain a MN. This procedure usually leads to a PGM with more parameters, that can lead to a longer training. Directed graphs, in which parent nodes do not share common children, are called moral. Consequently, BNs that are defined over moral graphs, can be converted to MNs without introducing additional edges and parameters.
Turning MNs into BNs is a more difficult task, both conceptually and computationally. First, the undirected graph needs to be triangulated, meaning that we introduce chords in cycles of $4$ or more and repeat this process until there are no such cycles left. Chords connect nodes, that are in the cycle, but are not already connected. This transformation usually leads to the introduction of a much larger number of edges than in the previous case. Finally, the undirected edges have to be turned into directed ones in an acyclic manner. Since chordal graphs are also moral, PGMs that are defined over chordal graphs represent a class of graphical models, that can be treated as either BNs or MNs and the conversion between the two is straightforward.
@@ II.1.4 Generative learning in PGMs
In MNs, the use of a normalizing constant couples the parameters across the whole network, which prevents us from decomposing the problem and estimating local groups of parameters separately. One of the computational ramifications of this global coupling is, that not even the parameter estimation with complete data can be solved in closed form (except for some special cases, that can essentially be reformulated as BNs [41]). This makes the use of iterative methods, such as gradient descent, unavoidable. Luckily, the objective function is convex, meaning that these methods are guaranteed to converge. However, they come with the disadvantage of having to run inference in each step to calculate the gradients and inference in MNs is #P-complete in general, which makes the distribution learning task formulated above fairly expensive or even intractable classically [42].
@@ II.2.1 General framework of QCBMs
QCBMs, introduced in Ref. [11], are paradigmatic generative QML models, that naturally inherit the Born rule and thus can be used to generate tunable and discrete probability distributions that approximate a target distribution. In this section, we review the general components of variational quantum algorithms in the context of QCBMs.
One of the key building blocks is the Ansatz. The Ansatz refers to a parametrized family of quantum circuits used as a hypothesis or approximation for solving the problem in question by iteratively optimizing the parameters to match the desired target behaviour or to minimize a cost function. We can differentiate problem-informed and problem-agnostic Ansätze.
Figure 3: General structure of the quantum circuit Ising Born machine, adapted from [13]. The circuit starts with Hadamard gates on each qubit, then a parametrized $n$-qubit unitary is applied with tunable parameters $\boldsymbol{\alpha}$, finally, a general parametrized $U_{f}$ gate acts on each qubit separately. The measurement is taken in the computational basis.
@@
$U_{f}(\mathbf{\Gamma},\mathbf{\Delta},\mathbf{\Sigma})=\exp{\left(i\sum\limits_{k=1}^{n}(\Gamma_{k}X_{k}+\Delta_{k}Y_{k}+\Sigma_{k}Z_{k})\right)}.$ ) . (4)
@@
There are multiple ways to characterize the distance of two distributions. The TV distance, shown in (2), is a good benchmark, but it is not feasible as a cost function, as it uses the probabilities explicitly. While in principle, one can use an approximation of the probability distributions based on a finite number of samples, the required number is exponential in general. One of the most common metrics used as a cost function in generative modelling is the Kullback-Leibler (KL) divergence
@@
Unfortunately, this function is also unsuitable for data-driven training, as it requires a large number of training samples, and it was shown not to be trainable for larger-scale QCBMs [43]. In Ref. [12] a more efficient cost function was introduced for gradient-based training of QCBMs. The idea was to compare the distance of the samples drawn from the target and the model distribution in a kernel feature space. This loss function is called the squared maximum mean discrepancy (MMD):
@@
Having a family of parametrized circuits and a cost function to characterize the distance between the circuit output and the target distribution, the final building block is the optimizer, that adjusts the circuit parameters given the cost function. Both gradient-based and gradient-free optimizers have been used to train QCBMs [12, 44]. In our numerical experiments, we only consider gradient-based optimization methods, but since our approach concentrates on Ansatz-design, it is compatible with other optimizers as well.
@@ II.2.2 Existing quantum circuit adaptations of PGMs
Liu and Wang [12] made the first explicit connection between the paradigmatic QCBM model and PGMs. They propose a framework, in which they first construct the Chow-Liu tree of a dataset [45] based on the mutual information between all pairs of bits for training samples. Having this tree graph, they propose a QCBM Ansatz, in which, the connectivity pattern of the $CNOT$ gates respects the graph structure. The Chow-Liu tree offers an effective approach for creating a second-order product approximation of a joint probability distribution. The corresponding graph represents a BN, that can also be regarded as a pairwise MN, but being a second-order approximation, it fails to detect higher-order correlations. Another explicit connection between BNs and QCBMs was formulated in Ref. [46], where the authors proposed a framework, that utilizes QCBMs for variational inference in PGMs.
@@
Besides the variational framework of general-purpose QCBMs, there have been several attempts to exactly implement PGMs on a quantum computer. BNs have an equivalent formulation in the computational basis measurements of a class of quantum circuits known as Bayesian quantum circuits (BQCs) [34]. These are defined such that the probability distribution they sample from, by measuring the given qubits in the computational basis, corresponds to the distribution defined by the corresponding BN. BQCs are implemented with uniformly controlled gates, that can be decomposed into one-qubit rotations and $CNOT$ gates [35, 47]. By definition, these circuits obey certain rules in accordance with the directed, acyclic nature of the underlying graph.
@@
In Ref. [36] the authors introduced a minimal extension to BQCs and presented unconditional proof of separation in the expressive power of BNs and the corresponding basis-enhanced BQCs (BBQCs). They showed that by letting the final measurement be in any local basis, this separation appears, that can be associated with quantum nonlocality and contextuality. They also pointed out, that both BQCs and their basis-enhanced versions can be efficiently simulated with classical tensor network methods, when the graphs are sparse enough.
@@
The literature on the quantum circuit implementation of Markov networks is more scarce. In Ref. [48] the authors identified a novel embedding technique of MNs into unitary operators that relies on their log-linear representation. They construct a Hamiltonian composed of Pauli-$Z$ terms and give a quantum algorithm, that implements the exponential of this Hamiltonian, meaning, that measuring the output qubits of the corresponding quantum circuit is equivalent to sampling the corresponding MN. The circuit, that implements this exponentiation, uses a special point-wise polynomial approximation and real part extraction, that might fail. For this reason, they have to measure ancillary qubits in order to determine whether this extraction was successful and potentially start everything over. Consequently, the success probability decreases exponentially with the number of maximal cliques. This can of course be amplified with quantum singular value transformation [49], but that further increases the required resources. Since Boltzmann machines form a subclass of pairwise MNs, their quantum circuit implementations can also be regarded as an adaptation of PGMs [17]. However, these models are quite restricted compared to general MNs, since they only consider pairwise correlations, usually in a bipartite manner.
@@ III Quantum Circuit Markov Random Fields
In this section, we present our results, starting with the definition of our QML model, proposed for generative learning in MNs. We then introduce our novel benchmark proposal and compare our model to both problem-agnostic QCIBMs and BBQCs through a series of numerical experiments. As a preliminary analysis of trainability, we investigate the scaling of the cost function variance for different types of graphs. Finally, we present our argument for a potential quantum advantage of our MN-based model class.